• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin Huang

    Log entry time 21:46:04 on November 04, 2008

    Entry number 246739


    keyword=Vertical Pumping Laser Fiber Status

    Vertical Pumping Laser Fiber is connected to Duo lasers and tested. -


    1. 4 vertical fibers (N8 N10 O4 O5) is connected to duo laser. We are able to power up to 4x25W. Fiber Temperature and laser spot is checked.
    2. Duo laser 1 (in middle of the rack), channel 1 can not reach a output power above 0.018W. It's not being used.
    3. From EPICS control console, all duo laser communication is showed to be bad. Faps are OK.

    Cable Map

    Index On EPICS Relatice Layout Location COM Port ID Port 1 Fiber Port 2 Fiber
    Duo 1 Middle 2E None None
    Duo 2 Bottom 1F N8 N10
    Duo 3 Upper 1A O4 O5

    A copy of this log entry has been emailed to: yqiang, chiran, jpchen